# MoEMeta: Mixture-of-Experts Meta Learning for Few-Shot Relational Learning

### v1

@@ 6 Conclusion
This paper presents MoEMeta, a novel meta-learning framework for few-shot relational learning in KGs. By integrating an MoE model to capture relational prototypes and a task-tailored adaptation mechanism to refine embeddings locally, MoEMeta significantly enhances the model’s ability to generalize across tasks while facilitating rapid adaptation tailored to individual tasks. Experimental results on three real-world KG datasets demonstrate that MoEMeta consistently outperforms state-of-the-art methods, showcasing its effectiveness for few-shot relation learning. In-depth ablation studies and visualizations further validate the importance of MoEMeta’s key components, confirming its efficacy in handling complex relational data in KGs with limited supervision.
@@ Appendix A Complexity Analysis
We analyze the computational complexity of our proposed MoEMeta for its key components. Table 5 presents the computational complexity of all components of MoEMeta, where $d$ denotes the dimension of entity embeddings, $n$ denotes the number of neighbors in attentive neighbor aggregation, and $K$ denotes the number of support triplets per task. For MoE, $M$ is the total number of experts, while $N$ is the number of activated experts in every forward pass.
As shown in Table 5, the complexity of MoEMeta scales linearly with the number of neighbors during the aggregation stage ($\mathcal{O}(Knd^{2})$) and linearly with the number of support triplets during MoE-based meta-knowledge learning ($\mathcal{O}(NKd^{2})$). The task-tailored local adaptation introduces only a minimal overhead of ($\mathcal{O}(d)$) in both parameters and computation. Overall, the computational cost is dominated by neighbor aggregation ($\mathcal{O}(Knd^{2})$), which is an essential component in current state-of-the-art methods. The additional computational cost and parameter overhead introduced by MoEMeta are negligible, making it efficient for few-shot KG relational learning.
Table 5: Complexity analysis of MoEMeta components with respect to the number of model parameters and the number of multiplication operations in each epoch. $d$ denotes the dimension of entity embeddings. $n$ denotes the number of neighbors per entity, and $K$ denotes the number of support triplets per task, where $K\ll n,d$. $M$ is the total number of experts, while $N$ is the number of activated experts in every forward pass.
@@ B.1 Reproducibility Details
Our method and all baselines are compared using the same evaluation protocol. The results of baseline methods are reproduced by using their official open-sourced implementations or adopted from their respective papers.
@@ B.2 Runtime Analysis
We conduct a runtime analysis comparing our proposed MoEMeta with other baselines. For fair comparison, we evaluate MoEMeta on a RTX3090 GPU to keep consistency with baseline runtimes sourced from the RelAdapter (Ran et al., 2024) paper. The total runtime for meta-training and the inference time per triplet during meta-testing are reported in seconds in Table 6.
@@
Appendix C Detailed Related Work
@@ C.1 Supervised KG Embedding Learning
KG embedding methods aim to learn a low-dimensional embedding space for entities and relations in KGs using input triplets as supervision. Two main categories of KG embedding methods for KG relational learning include: (1) Translation-based methods, such as TransE (Bordes et al., 2013), TransH (Wang et al., 2014), TransD (Ji et al., 2015), and TransR (Lin et al., 2015), which assumes a translational relationship between entities and relations for learning a unified embedding space. (2) Semantic matching methods, such as ComplEx (Trouillon et al., 2016), RotatE (Sun et al., 2019), and ComplEx-N3 (Lacroix et al., 2018), which improve the modeling of relation patterns in a vector/complex space. However, these methods rely on a large amount of triplets for training. Their performance deteriorates at few-shot settings, where only a handful of training triplets are available for each relation.
@@ C.2 Few-Shot Learning and Meta Learning
Few-shot learning aims to train a supervised model that can generalize to new tasks with only limited amounts of training data (Fei-Fei et al., 2006). Meta-learning (Finn et al., 2017; Santoro et al., 2016; Ravi and Larochelle, 2017; Nichol and Schulman, 2018; Ren et al., 2018) has emerged as a powerful paradigm for enabling rapid learning on few-shot tasks by leveraging meta-knowledge learned from meta-training tasks. Relative to metric-based approaches (Jake Snell and Zemel, 2017; Vinyals et al., 2016), optimization-based meta-learning methods have gained significant attention due to their strong generalization capacities. These methods cast meta-learning into a bi-level optimization problem; the inner loop fine-tunes a base model on individual tasks, and the outer loop optimizes the base model for quick adaptation across tasks. MAML (Finn et al., 2017) is the most representative algorithm in this category, which learns global parameters as an initialization to be shared among task-specific models for fast adaptation within one or a few steps of gradient descent. Due to its flexible formulation, MAML has been applied to various domains including vision, robotics, lifelong learning, and beyond. Subsequent variants of MAML (Finn et al., 2018; Von Oswald et al., 2021) improve upon its adaptability by optimizing model initialization or learning adaptive learning rates for fast convergence (Sun and Gao, 2023; Zou et al., 2021; Finn et al., 2018). Our work focuses on learning well-generalized meta-knowledge across tasks while incorporating local adaptation to improve few-shot relational learning in KGs.
@@ C.3 Few-Shot Relation Learning
Most mainstream FSRL methods are based on meta-learning (Finn et al., 2017), which focuses on learning an initial meta-prior from existing tasks that can generalize to new few-shot relations. MetaR (Chen et al., 2019) learns relation-specific meta-information for adaptation, by simply averaging the representations of all support triplets. GANA (Niu et al., 2021) integrates meta-learning with TransH (Wang et al., 2014) and devises a gated and attentive neighbor aggregator to address sparse neighbors for learning informative entity embeddings. HiRe (Wu et al., 2023) exploits multi-granularity relational information to learn meta-knowledge that better generalizes to unseen relations. RelAdapter (Ran et al., 2024) adopts a feedforward network with a residual layer to adapt relation-meta to the target relation. Additionally, pre-trained context information of related entities are provided to augment entity embeddings.
Despite great progress, the aforementioned meta-learning based methods suffer from two key limitations. First, these methods neglect common relational patterns among diverse relations when learning meta-knowledge during meta-training. Second, they fail to effectively incorporate local, task-specific information essential for fast adaptation, as global meta-knowledge may not be optimally tailored to the unique characteristics of individual tasks. Our work is proposed to address these critical gaps within the meta-learning framework, enabling more effective model generalization and adaptation under few-shot settings.
Our work is also related to recent graph-based and contrastive meta-learning methods such as (Liu et al., 2022) and (Liu et al., 2024). While these approaches are relevant to few-shot learning and meta-learning on node-level tasks for homogeneous graphs, they are not directly applicable to our setting, which focuses on few-shot relational learning in KGs—that centers on relation-level tasks. In contrast, our method, MoEMeta, specifically targets the key challenges of few-shot relational learning, such as the complex and diverse relational patterns in KGs. This distinction sets our work apart from general graph few-shot learning or meta-learning methods that target node-level tasks.

### v2

@@ 6 Conclusion
This work presents MoEMeta, a novel meta-learning framework for few-shot relational learning in KGs. By integrating an MoE model to capture globally shared relational prototypes with a task-tailored adaptation mechanism to refine embeddings locally, MoEMeta effectively balances cross-task generalization and rapid task-specific adaptation. Extensive experiments on three real-world KG datasets demonstrate that MoEMeta consistently outperforms state-of-the-art methods, showcasing its effectiveness for few-shot relation learning. Comprehensive ablation studies and visualizations further validate the importance of MoEMeta’s key components, confirming its efficacy in modeling complex relational data in KGs under limited supervision.
Appendix
@@ Appendix A Complexity Analysis
In this section, we analyze the computational complexity of our proposed MoEMeta for its key components. Table 5 summarizes the computational complexity of each component, where $d$ denotes the dimension of entity embeddings, $n$ denotes the number of neighbors in attentive neighbor aggregation, and $K$ denotes the number of support triplets per task. For MoE-based meta-knowledge learning, $M$ indicates the total number of experts, and $N$ indicates the number of experts activated in each forward pass.
As shown in Table 5, the complexity of MoEMeta scales linearly with the number of neighbors per entity during the aggregation stage ($\mathcal{O}(Knd^{2})$) and linearly with the number of support triplets during MoE-based meta-knowledge learning ($\mathcal{O}(NKd^{2})$). The task-tailored local adaptation introduces only a minimal overhead of ($\mathcal{O}(d)$) in both parameters and computation. Overall, the computational cost is dominated by neighbor aggregation ($\mathcal{O}(Knd^{2})$), which is an essential component in current state-of-the-art FSRL methods. The additional computational and parameter overhead introduced by MoEMeta is negligible, making it efficient for FSRL in KGs.
Table 5: Complexity analysis of MoEMeta components with respect to the number of model parameters and the number of multiplication operations in each epoch. $d$ denotes the dimension of entity embeddings. $n$ denotes the number of neighbors per entity, and $K$ denotes the number of support triplets per task, where $K\ll n,d$. $M$ is the total number of experts, while NN is the number of experts activated in each forward pass.
@@ B.1 Reproducibility Details
In our experiments, our method and all baselines are compared using the same evaluation protocol. The results of baseline methods are reproduced by using their official open-sourced implementations or adopted from their respective papers.
@@ B.2 Runtime Analysis
We also perform a runtime analysis comparing MoEMeta with other baselines. For a fair comparison, we evaluate MoEMeta on an RTX3090 GPU to keep consistency with baseline runtimes reported in RelAdapter (Ran et al., 2024). Table 6 reports the total runtime of meta-training and the inference time per triplet during meta-testing in seconds.
@@
Appendix C Detailed and Additional Related Work
For a more comprehensive discussion of related literature beyond the main paper, we provide additional related work in this section.
@@ C.1 Supervised KG Embedding Learning
KG embedding methods aim to learn a low-dimensional embedding space for entities and relations in KGs using input triplets as supervision. They broadly fall into two categories for KG relational learning: (1) Translation-based methods, such as TransE (Bordes et al., 2013), TransH (Wang et al., 2014), TransD (Ji et al., 2015), and TransR (Lin et al., 2015), which assume a translational relationship between entities and relations for learning a unified embedding space; and (2) Semantic matching methods, such as ComplEx (Trouillon et al., 2016), RotatE (Sun et al., 2019), and ComplEx-N3 (Lacroix et al., 2018), which improve the modeling of relational patterns in a vector or complex space. However, these methods rely on large amounts of triplets for training, and their performance degrades significantly under few-shot settings, where only a handful of training triplets are available for each relation.
@@ C.2 Few-Shot Learning and Meta Learning
Few-shot learning aims to train a supervised model that can generalize to new tasks with only limited amounts of training data (Fei-Fei et al., 2006). Meta-learning (Finn et al., 2017; Santoro et al., 2016; Ravi and Larochelle, 2017; Nichol and Schulman, 2018; Ren et al., 2018) has emerged as a powerful paradigm for enabling rapid learning on few-shot tasks by leveraging meta-knowledge learned from meta-training tasks. Relative to metric-based approaches (Jake Snell and Zemel, 2017; Vinyals et al., 2016), optimization-based meta-learning methods have gained significant attention due to their strong generalization capabilities. These methods cast meta-learning into a bi-level optimization problem; the inner loop fine-tunes a base model on individual tasks, and the outer loop optimizes the base model for quick adaptation across tasks. MAML (Finn et al., 2017) is the most representative algorithm in this category, which learns global parameters as an initialization to be shared among task-specific models for fast adaptation using one or a few steps of gradient descent updates. Due to its flexible formulation, MAML has been applied to various domains, including computer vision, robotics, lifelong learning, and beyond. Subsequent variants of MAML (Finn et al., 2018; Von Oswald et al., 2021) improve upon its adaptability by optimizing model initialization or learning adaptive learning rates for faster convergence (Sun and Gao, 2023; Zou et al., 2021; Finn et al., 2018). Our work focuses on learning well-generalized meta-knowledge across tasks while integrating task-specific local adaptation to improve few-shot relational learning in KGs.
@@ C.3 Few-Shot Relation Learning
Most mainstream FSRL methods are based on meta-learning (Finn et al., 2017), which focuses on learning an initial meta-prior from existing tasks that can generalize to new few-shot relations. MetaR (Chen et al., 2019) learns relation-specific meta-information for adaptation, by simply averaging the representations of all support triplets. GANA (Niu et al., 2021) integrates meta-learning with TransH (Wang et al., 2014) and devises a gated and attentive neighbor aggregator to address sparse neighbors for learning informative entity embeddings. HiRe (Wu et al., 2023) exploits multi-granularity relational information to learn meta-knowledge that better generalizes to unseen relations. RelAdapter (Ran et al., 2024) adopts a shallow feedforward network with a residual connection to adapt relation-meta within each target task. Additionally, pre-trained context information of related entities is used to augment entity embeddings.
Despite their great progress, the aforementioned meta-learning based methods suffer from two key limitations. First, these methods neglect common relational patterns shared across tasks when learning meta-knowledge during meta-training. Second, they fail to effectively incorporate local, task-specific information essential for fast adaptation, as global meta-knowledge may not be optimally tailored to the unique characteristics of individual tasks. Our work is thus proposed to address these critical gaps within the meta-learning framework, enabling more effective model generalization and adaptation under few-shot settings.
Our work is also related to recent graph-based and contrastive meta-learning methods such as (Liu et al., 2022) and (Liu et al., 2024). While these approaches are relevant to few-shot learning and meta-learning on node-level tasks for homogeneous graphs, they are not directly applicable to our setting, which focuses on few-shot relational learning in KGs—that centers on relation-level tasks. Our method, MoEMeta, specifically targets the key challenges of few-shot relational learning, such as the complex and diverse relational patterns in KGs. This distinction sets our work apart from general graph few-shot learning or meta-learning methods that target node-level tasks.